# Product Digital Twin Supporting End-of-life Phase of Electric Vehicle Batteries Utilizing Product-Process-Resource Asset Network

### v1

@@ V Evaluation and Discussion
We have already used the PAN paradigm in the frame of other research projects and we found it to be efficient for capturing expert knowledge relevant for multi-disciplinary mechatronic production systems. However, the standard PAN does not provide support for remanufacturing or recycling processes. Therefore, we have extended the PAN paradigm in this paper and we have addressed the research question RQ1. The proposed extension can express bi-directional manufacturing operations seamlessly in one graph, applicable for (i) for manufacturing (i.e., forward direction through the Bi-PAN) and (ii) for remanufacturing/disassembling (i.e., backward direction through the Bi-PAN).
Addressing the research question RQ-2, the important benefit of the PAN paradigm (incl. the original one and the extended version proposed in this paper) is the compliance with industrial standards such as ISA-95 or AutomationML. It can be used in the conjunction with model-checking techniques to check the consistency of the overall PAN. The very same model in the Bi-PAN can be utilized for manufacturing and remanufacturing in intelligent factories. Added distinction between elementary/atomic products and (composite) products in PAN improves readability for humans but it is also useful for automated processing of PAN graphs.
@@
Comparing and evaluating the proposed Bi-PAN approach for the digital twin design with the traditional DT design techniques is evaluated with Likert’s scale in Fig. 5.
The conventional design techniques are suitable for their original purpose only, it means for manufacturing/product assembling (see ++ in the upper left-hand side corner of Fig. 5). On the other hand, conventional design methods for DT do not provide enough support for product disassembling or remanufacturing, which are typically out-of-scope of these techniques (or that require a separate brand-new DT tailored for these purposes). The proposed Bi-PAN supports both assembling and disassembling of products. A trade-off for supporting both is a slightly more complex model, requiring more modeling effort and time (approx. 10–20%). This is the reason why the product assembling based on the Bi-PAN is evaluated with (+), compared to (++) in the case of the conventional DT design. The significant benefit of the Bi-PAN is a support for remanufacturing or reparation of products, as it is shown in the EV battery use-case in the previous section (++ on the Likert’s scale in Fig. 5). With conventional DT design, support for disassembling is missing and the combination of product assembling and disassembling in the remanufacturing processes is not possible at all.
@@ VI Conclusion and Future Work
Remanufacturing and recycling are key phases of the EV battery life-cycle in order to provide economic and ecologic sustainability of their usage. In this paper, the approach is based on the use of the product digital twin technology to persistently keep knowledge about the whole battery, their modules, but also process steps to assemble, disassemble, or repair the battery. The proposed approach for the product digital twin design is based on the proposed extension of product-process-resource asset network called Bi-PAN.
The PDT based on Bi-PAN is used at the End-of-life phase of the battery. It enables to query required process to disassemble or repair the EV battery that can be done either manually or on a flexible production system providing the required skills.
@@
In the future work, we will focus on a better integration of the proposed PAN-based approach with the Asset Administration Shell for EV batteries. We would also like to create a PAN model for a selected EV battery in full details and discuss it with the car vendor of the EV, together with elaborating on the scalability and potential limitations of this approach in the real-life scale.
@@ Acknowledgment
This work was supported by the Grant Agency of the Czech Technical University in Prague, grant No. SGS24/125/OHK2/3T/12 and the project “Regeneration of used batteries from Electric Vehicles” (Slovak ITMS2014+ code 313012BUN5), the project is part of the Important Project of Common European Interest (IPCEI) called the European Battery Innovation (code OPII-MH/DP/2021/9.5-34), which was announced as a part of Operational Program Integrated Infrastructure (EZOP ID 71235). This work was co-funded by the European Union under the project Robotics and advanced industrial production – ROBOPROX (reg. no. CZ.02.01.01/00/22_008/0004590).

### v2

@@ V Evaluation and Discussion
We have already used the (standard) PAN paradigm in the frame of other research projects and we found it to be efficient for capturing expert knowledge relevant for multi-disciplinary mechatronic production systems. However, the standard PAN does not provide support for remanufacturing or recycling processes. For this reason, we have extended the PAN paradigm in this paper and we have thus addressed the research question RQ1. The proposed extension can express bi-directional manufacturing operations seamlessly in one graph, applicable (i) for manufacturing (i.e., forward direction through the Bi-PAN), and (ii) for remanufacturing/disassembling (i.e., backward direction through the Bi-PAN).
Addressing the research question RQ-2, the important benefit of the PAN paradigm (incl. the original one and the extended version proposed in this paper) is compliance with industrial standards such as ISA-95 or AutomationML. It can be used in conjunction with model-checking techniques to check the consistency of the overall PAN. The very same model in the Bi-PAN can be utilized for manufacturing and remanufacturing in intelligent factories. Added distinction between elementary/atomic products and (composite) products in PAN improves readability for humans but it is also useful for automated processing of PAN graphs.
@@
Comparing and evaluating the proposed Bi-PAN approach for the digital twin design with the traditional DT design techniques is evaluated with Likert’s scale in Fig. 4. The conventional design techniques are suitable for their original purpose only, it means for manufacturing/product assembling (see “++” in the upper left-hand side corner of Fig. 4). On the other hand, conventional design methods for DT do not provide enough support for product disassembling or remanufacturing, which are typically out-of-scope of these techniques (or that require a separate brand-new DT tailored for these purposes), see “–” in product disassembling and remanufacturing. The proposed Bi-PAN supports both assembling and disassembling of products (see “+” and “++” in the right-hand side column). A trade-off for supporting both is a slightly more complex model, requiring more modeling effort and time. This is the reason why the product assembling based on the Bi-PAN is evaluated with “+”, compared to “++” in the case of the conventional DT design. The significant benefit of the Bi-PAN is the support for remanufacturing or reparation of products, as it is shown in the EV battery use-case in the previous section (“++” on the Likert’s scale in Fig. 4). With conventional DT design, support for disassembling is missing and the combination of product assembling and disassembling in the remanufacturing processes is not possible.
@@ VI Conclusion and Future Work
Remanufacturing and recycling are key phases of the EV battery life-cycle in order to provide economic and ecologic sustainability for their usage. The approach proposed in this paper utilizes the product digital twin technology to persistently keep knowledge about the whole battery, and its modules, but also process steps to assemble, disassemble, or repair the battery. The proposed approach for the product digital twin design is based on the proposed extension of the product-process-resource asset network called Bi-PAN.
The PDT based on Bi-PAN is used at the end-of-life phase of the battery. It enables querying of particular required processes including disassembling or repairing the EV battery which can be done either manually or on a flexible production system providing the required skills.
@@
In the future work, we are going to focus on a better integration of the proposed PAN-based approach with the Asset Administration Shell for EV batteries. We would also like to create a PAN model for a selected EV battery in full detail and discuss it with the car vendor of the EV, together with elaborating on the scalability and potential limitations of this approach on a real-life scale.
@@ Acknowledgment
This paper has been supported by the Grant Agency of the Czech Technical University in Prague, under grant No. SGS24/125/OHK2/3T/12 as well as by the project “Regeneration of used batteries from Electric Vehicles” (i.e., Slovak ITMS2014+ code 313012BUN5), the project is a part of the “Important Project of Common European Interest” (IPCEI) called the “European Battery Innovation” (i.e., code OPII-MH/DP/2021/9.5-34), which is a part of the “Operational Program Integrated Infrastructure” (EZOP No. 71235). The work was co-funded by the European Union (EU) under the project Robotics and advanced industrial production – ROBOPROX (reg. no. CZ.02.01.01/00/22_008/0004590).